# Class-Aware PillarMix: Can Mixed Sample Data Augmentation Enhance 3D Object Detection with Radar Point Clouds?

### v2

@@ Abstract
Due to the significant effort required for data collection and annotation in 3D perception tasks, mixed sample data augmentation (MSDA) has been widely studied to generate diverse training samples by mixing existing data. Among these methods, MixUp is a prominent approach that generates new samples by linearly combining two existing ones, using a mix ratio sampled from a $\beta$ distribution. This simple yet powerful method has inspired numerous variations and applications in 2D and 3D data domains. Recently, many MSDA techniques have been developed for point clouds, but they mainly target LiDAR data, leaving their application to radar point clouds largely unexplored. In this paper, we examine the feasibility of applying existing MSDA methods to radar point clouds and identify several challenges in adapting these techniques. These obstacles stem from the radar’s irregular angular distribution, deviations from a single-sensor polar layout in multi-radar setups, and point sparsity. To address these issues, we propose Class-Aware PillarMix (CAPMix), a novel MSDA approach that applies MixUp at the pillar level in 3D point clouds, guided by class labels. Unlike methods that rely a single mix ratio to the entire sample, CAPMix assigns an independent ratio to each pillar, boosting sample diversity. To account for the density of different classes, we use class-specific distributions: for dense objects (e.g., large vehicles), we skew ratios to favor points from another sample, while for sparse objects (e.g., pedestrians), we sample more points from the original. This class-aware mixing retains critical details and enriches each sample with new information, ultimately generating more diverse training data. Experimental results demonstrate that our method not only significantly boosts performance but also outperforms existing MSDA approaches across two datasets (Bosch Street and K-Radar). We believe that this straightforward yet effective approach will spark further investigation into MSDA techniques for radar data.
@@ I INTRODUCTION
Autonomous driving technologies have advanced rapidly, driven by significant progress in deep learning. Yet, 3D perception tasks remain highly data-dependent, typically requiring extensively annotated datasets for reliable performance. Collecting and labeling such data in real-world driving scenarios can be highly time-consuming and labor-intensive. To tackle these challenges, mixed sample data augmentation (MSDA) is widely studied [1, 2, 3, 4, 5]. By blending LiDAR point clouds from different samples, MSDA effectively expands and diversifies training sets, resulting in notable gains in fully supervised, semi-supervised, and domain adaptation contexts, particularly when data and/or labels are scarce.
555510101010252525255050505035353535404040404545454550505050100% Data ResultData Portion (%)$\text{mAP}_{\text{BEV}}\text{ (\%)}$%)w/o MSDACAPMix
@@
Radar sensors transmit electromagnetic waves to measure the range, angle, and velocity of surrounding targets. Their robustness in adverse weather conditions and low cost made them increasingly important for 3D perception tasks in autonomous driving [6]. Compared to LiDARs, radar point clouds are typically sparser and noisier, which intensifies the data scarcity problem. Although recent research has focused on designing advanced models to improve radar perception [7, 8, 9, 10], their performance in low-data scenarios remains relatively limited and largely unexplored. Since both radar and LiDAR point clouds share a similar format (3D coordinates plus additional attributes), we hypothesize that MSDA methods are also well-suited for radar 3D perception tasks. To investigate this, we assess how state-of-the-art MSDA techniques, originally designed for LiDAR, translate to radar data. Our findings reveal several obstacles, including the irregular angular distribution from radar’s radiation pattern, inherent point sparsity, and deviations from a single-sensor polar distribution in multi-radar setups, all of which limit the direct application of existing MSDA approaches.
In this paper, we propose Class-Aware PillarMix (CAPMix), a novel approach that mixes radar point clouds with independently determined mix ratios based on class information: We adopt pillars (columnar grid cells in 3D space) as the operational units for MSDA, as they provide an effective representation of the 3D space. Unlike PillarMix [5], which directly swaps points, we apply MixUp [11] within the pillars to achieve class-aware data augmentation. Similar to MixUp [11], our method samples a mix ratio from a specified $\beta$ distribution to determine the proportion of points taken from the two inputs. To improve sample diversity, each pillar is assigned its own mix ratio. Leveraging class information from the labels, we apply distinct $\beta$ distributions to different pillars for ratio sampling. This strategy ensures that the generated samples are augmented while retaining crucial object-level details.
@@
  • We validate CAPMix on two radar datasets (Bosch Street [12] and K-Radar [13]) of different scales, showcasing its superior performance over other state-of-the-art techniques in terms of $AP_{BEV}$ and $AP_{3D}$.
@@ II Related Works
MSDA generates new samples by mixing the existing training samples. It was first introduced for image tasks in MixUp [11], which combines input pixels and output labels in a specified ratio. Consequently, it has become an effective technique to enhance model robustness and is compatible with traditional augmentations, such as random rotation and scaling, while also generating more diverse samples for training. Building on this approach, CutMix [14] randomly crops a rectangular region from one image and pastes it onto another. More recently, PuzzleMix [15] incorporates saliency information from the backward gradient into the mixing process, while AutoMix [16] trains a learnable MixUp generator under the supervision of mixed labels. These pixel-level mixing methods have attracted increasing attention and demonstrate both effectiveness and generalizability in 2D image perception tasks [17]. However, fundamental differences between 2D images and 3D point clouds prevent most of these approaches from being directly applied to 3D data.
Recent works have extended MSDA to 3D perception tasks [2, 5, 4, 3] to enhance 3D perception performance across various tasks, such as object detection and semantic segmentation. Given two point clouds $s^{A}$, $s^{B}$ with their labels $y^{A}$ and $y^{B}$, these methods select points from each of the samples to generate a new sample $\tilde{s}$ and its corresponding label $\tilde{y}$. For example, MixUp [3] randomly selects points by a given mix ratio. PolarMix [4] selects one area from each of the two point clouds based on azimuth angle. Similarly, LaserMix [2] partitions the point clouds into multiple regions based on elevation angle, then merges these regions from both point clouds in an alternating manner. Furthermore, PillarMix [5] performs alternating selection on pillars instead of angles.
@@
    MixUp [11] was adapted for LiDAR point clouds in [3]. Binary masks $K$ and $Q$ are employed to select the points such that $\frac{|K=1|}{|K|}=\lambda$ and $\frac{|Q=1|}{|Q|}=1-\lambda$, where $\lambda$ is sampled from a given distribution $\beta$. The output is computed as $\tilde{s}=(K\odot s^{A})\oplus(Q\odot s^{B})$ and the new label $\tilde{y}=y^{A}\oplus y^{B}$. Here, $\odot$ and $\oplus$ represent masking operation and concatenation, respectively. A common choice of $\beta$ distribution is $\beta(2,2)$, which concentrates most values around 0.5. However, radar point clouds can be extremely sparse in certain regions, meaning that applying a single mix ratio uniformly across all points may be suboptimal and limit diversity.
@@
    PolarMix [4] randomly divides the field of view into two azimuth regions and swaps all points within the selected angular interval. As shown in Fig. 3(a), a random 90 region is chosen for this exchange, with all points and corresponding labels being swapped. Additionally, points within bounding boxes in $s^{B}$ are copied and rotated around the origin by a randomly selected angle within a predefined range into $\tilde{s}$ to enhance object diversity. Unlike LiDAR sensors, which offer uniform 360 detection, radar point clouds encounter irregularities in their angular distribution due to the radiation pattern, as shown in Fig. 2(a). This leads to significant variations in detection capability across the main lobe and side lobes, especially along the azimuth axis. Accordingly, substantial rotation operations such as PolarMix will disrupt the inherent angular distribution of the point cloud.
@@
    LaserMix [2] exploits the inclination angle for partitioning by leveraging the distinctive pattern distribution of laser beams. As explained in Fig. 3(b), the points clouds are divided into $N$ equally spaced segments based on the elevation angular interval. An alternating selection of point segments from $s^{A},s^{B}$ is then utilized to construct $\tilde{s}$. However, multi-radar setups are frequently adopted to enhance radar perception, for instance in the Bosch Street dataset [12], which employs five front-facing radars. As Fig. 2(b) shows, these multiple sensors emit beams from different coordinate systems, so the spatial priors established by laser partitions mentioned in LaserMix [2] no longer hold. Additionally, radar sensors typically have a lower resolution in the elevation angle and a narrower vertical field of view compared to LiDAR [18] (Fig. 2(c)). Thus, mixing based on elevation angle or polar coordinates may not be applicable to radar sensors due to their limited elevation performance and the effects of multi-sensor aggregation.
@@
    PillarMix [5] directly switches the points and labels within selected pillars to generate mixed samples Compared to LiDAR point clouds, radar point clouds are significantly sparser. Replacing portions of these sparse objects with irrelevant or empty pillars may remove critical information needed for accurate predictions. Consequently, training on such augmented data can degrade performance, making PillarMix less effective than MixUp for these objects. Hence, directly switching entire pillars may be suboptimal for radar point clouds, especially when dealing with sparse objects.
As discussed above, the distinct point distribution patterns in radar point clouds, combined with their inherent sparsity and the use of multi-radar sensor fusion, differ significantly from 360 LiDAR data. Therefore, aforementioned methods developed for LiDAR data often fall short when applied to radar. Later results will reveal that some of these approaches not only underperform compared to general mixed-sample techniques like MixUp [11] but can even negatively impacting perception performance.
@@ III Approach
Figure 4: Overall framework comparison of PillarMix and our proposed CAPMix. Both methods share the pillarization step, but PillarMix simply switches pillars between two samples $s^{A},s^{B}$. In contrast, CAPMix randomly selects pillars and then performs class-aware MixUp using different $\beta$ distributions ${\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}},{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}},{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$ based on the classes in the pillars.
@@ III-A Overview
Inspired by PillarMix [5], we propose CAPMix, a class-aware variant that accounts for the sparsity of radar point clouds. The overall methods of both PillarMix and CAPMix are illustrated in Fig. 4. Like PillarMix, CAPMix begins by pillarizing two point clouds $s^{A},s^{B}$ using a specified pillar size $\phi$ (covering an area of $\phi\times\phi$ in Bird’s-Eye View). However, instead of using non-adjacent cross-selection as in PillarMix, all the pillars $P$ are randomly divided into $[P_{c},P_{u}]$ based on a ratio $\alpha$. Next, MixUp is applied to each pillar, with the mix ratio independently sampled from a predefined $\beta$ distribution. For $P_{c}$, the mix ratio is class-aware, with pillars assigned a mix ratio from ${\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}},\,{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}},\,\textrm{or}\,\,{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$ based on their object density (sparse, moderate, or dense). To enhance diversity while preserving original information, ${\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}}$ is also applied to all pillars in $P_{u}$ without considering class labels. As a data augmentation technique, CAPMix is employed exclusively during training, while the inference stage remains identical to the standard process.
@@ III-C Class-aware $\beta$ distributions in $P_{c}$
1:Points and labels of two radar samples: $\{s^{A},\,y^{A}\},\{s^{B},\,y^{B}\}$; Pillar Size $\phi$; Pillar select ratio $\alpha$; MSDA probability $\delta$; MixUp $\beta$ distribution ${\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}},\,{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}},\,{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$ for $cls_{moderate},cls_{sparse},cls_{dense}$.
@@
6:    $[P_{c},\,P_{u}]=\textrm{RandomSelect}([P_{A},\,P_{B}],\,\alpha),$
@@
To ensure the model retains sufficient information for accurate predictions, the class-aware mix ratio is applied only to $P_{c}\,$, rather than all pillars. Applying it universally would result in excessive information loss, particularly for large objects, making them significantly harder for the model to detect. For instance, if ${\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$ were applied to dense objects in $P_{u}$, every part of the object would receive a $\lambda$ value close to 0, leading to the removal of most points. In contrast, using ${\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}}$ for $P_{u}$ ensures that enough points are preserved, maintaining the integrity of the object and facilitating reliable predictions. Compared to PillarMix [5], which retains all points in $P_{u}\,$, applying a moderate mix ratio provides a balanced approach, preserving sufficient information for effective detection while still benefiting from augmentation. The operation can be mathematically expressed as follows:
@@
$\tilde{s}$ =$=\{\tilde{p}_{c}\}\oplus\{\tilde{p}_{u}\},$
@@
Our method leverages bounding box labels while operating on radar data pillars, recognizing that radar points can extend beyond the bounding box, particularly for moving objects  [19]. By applying pillar-wise operations, we ensure that these points, along with non-object points, are included in the mixing process. Moreover, we use object labels rather than sparsity to distinguish pillars, ensuring sparse objects are not overshadowed or erased when sharing a pillar with dense objects. This careful handling of sparse points with high priority is crucial for maintaining data integrity.

### v3

@@ Abstract
Due to the significant effort required for data collection and annotation in 3D perception tasks, mixed sample data augmentation (MSDA) has been widely studied to generate diverse training samples by mixing existing data. Among these methods, MixUp is a prominent approach that generates new samples by linearly combining two existing ones, using a mix ratio sampled from a $\beta$ distribution. This simple yet powerful method has inspired numerous variations and applications in 2D and 3D data domains. Recently, many MSDA techniques have been developed for point clouds, but they mainly target LiDAR data, leaving their application to radar point clouds largely unexplored. In this paper, we examine the feasibility of applying existing MSDA methods to radar point clouds and identify several challenges in adapting these techniques. These obstacles stem from the radar’s irregular angular distribution, deviations from a single-sensor polar layout in multi-radar setups, and point sparsity. To address these issues, we propose Class-Aware PillarMix (CAPMix), a novel MSDA approach that applies MixUp at the pillar level in 3D point clouds, guided by class labels. Unlike methods that rely a single mix ratio to the entire sample, CAPMix assigns an independent ratio to each pillar, boosting sample diversity. To account for the density of different classes, we use class-specific distributions: for dense objects (e.g., large vehicles), we skew ratios to favor points from another sample, while for sparse objects (e.g., pedestrians), we sample more points from the original. This class-aware mixing retains critical details and enriches each sample with new information, ultimately generating more diverse training data. Experimental results demonstrate that our method not only significantly boosts performance but also outperforms existing MSDA approaches across two datasets (Bosch Street and K-Radar). We believe that this straightforward yet effective approach will spark further investigation into MSDA techniques for radar data. Code is available at https://github.com/boschresearch/CAPMIX.
@@ I INTRODUCTION
Autonomous driving technologies have advanced rapidly, driven by significant progress in deep learning. Yet, 3D perception tasks remain highly data-dependent, typically requiring extensively annotated datasets for reliable performance. Collecting and labeling such data in real-world driving scenarios can be highly time-consuming and labor-intensive. To tackle these challenges, mixed sample data augmentation (MSDA) is widely studied [park2022unified, kong2023lasermix, wang2023ssda3d, xiao2022polarmix, wu2024pillarmix]. By blending LiDAR point clouds from different samples, MSDA effectively expands and diversifies training sets, resulting in notable gains in fully supervised, semi-supervised, and domain adaptation contexts, particularly when data and/or labels are scarce.
[…]
@@
Radar sensors transmit electromagnetic waves to measure the range, angle, and velocity of surrounding targets. Their robustness in adverse weather conditions and low cost made them increasingly important for 3D perception tasks in autonomous driving [srivastav2023radars]. Compared to LiDARs, radar point clouds are typically sparser and noisier, which intensifies the data scarcity problem. Although recent research has focused on designing advanced models to improve radar perception [liu2023smurf, ulrich2022improved, musiat2024radarpillars, palffy2022vod], their performance in low-data scenarios remains relatively limited and largely unexplored. Since both radar and LiDAR point clouds share a similar format (3D coordinates plus additional attributes), we hypothesize that MSDA methods are also well-suited for radar 3D perception tasks. To investigate this, we assess how state-of-the-art MSDA techniques, originally designed for LiDAR, translate to radar data. Our findings reveal several obstacles, including the irregular angular distribution from radar’s radiation pattern, inherent point sparsity, and deviations from a single-sensor polar distribution in multi-radar setups, all of which limit the direct application of existing MSDA approaches.
In this paper, we propose Class-Aware PillarMix (CAPMix), a novel approach that mixes radar point clouds with independently determined mix ratios based on class information: We adopt pillars (columnar grid cells in 3D space) as the operational units for MSDA, as they provide an effective representation of the 3D space. Unlike PillarMix [wu2024pillarmix], which directly swaps points, we apply MixUp [zhang2017mixup] within the pillars to achieve class-aware data augmentation. Similar to MixUp [zhang2017mixup], our method samples a mix ratio from a specified $\beta$ distribution to determine the proportion of points taken from the two inputs. To improve sample diversity, each pillar is assigned its own mix ratio. Leveraging class information from the labels, we apply distinct $\beta$ distributions to different pillars for ratio sampling. This strategy ensures that the generated samples are augmented while retaining crucial object-level details.
@@
  • We validate CAPMix on two radar datasets (Bosch Street [armanious2024bosch] and K-Radar [paek2022k]) of different scales, showcasing its superior performance over other state-of-the-art techniques in terms of $AP_{BEV}$ and $AP_{3D}$.
@@ II Related Works
MSDA generates new samples by mixing the existing training samples. It was first introduced for image tasks in MixUp [zhang2017mixup], which combines input pixels and output labels in a specified ratio. Consequently, it has become an effective technique to enhance model robustness and is compatible with traditional augmentations, such as random rotation and scaling, while also generating more diverse samples for training. Building on this approach, CutMix [yun2019cutmix] randomly crops a rectangular region from one image and pastes it onto another. More recently, PuzzleMix [kim2020puzzle] incorporates saliency information from the backward gradient into the mixing process, while AutoMix [liu2022automix] trains a learnable MixUp generator under the supervision of mixed labels. These pixel-level mixing methods have attracted increasing attention and demonstrate both effectiveness and generalizability in 2D image perception tasks [naveed2024mixsurvey]. However, fundamental differences between 2D images and 3D point clouds prevent most of these approaches from being directly applied to 3D data.
Recent works have extended MSDA to 3D perception tasks [kong2023lasermix, wu2024pillarmix, xiao2022polarmix, wang2023ssda3d] to enhance 3D perception performance across various tasks, such as object detection and semantic segmentation. Given two point clouds $s^{A}$, $s^{B}$ with their labels $y^{A}$ and $y^{B}$, these methods select points from each of the samples to generate a new sample $\tilde{s}$ and its corresponding label $\tilde{y}$. For example, MixUp [wang2023ssda3d] randomly selects points by a given mix ratio. PolarMix [xiao2022polarmix] selects one area from each of the two point clouds based on azimuth angle. Similarly, LaserMix [kong2023lasermix] partitions the point clouds into multiple regions based on elevation angle, then merges these regions from both point clouds in an alternating manner. Furthermore, PillarMix [wu2024pillarmix] performs alternating selection on pillars instead of angles.
@@
    MixUp [zhang2017mixup] was adapted for LiDAR point clouds in [wang2023ssda3d]. Binary masks $K$ and $Q$ are employed to select the points such that $\frac{|K=1|}{|K|}=\lambda$ and $\frac{|Q=1|}{|Q|}=1-\lambda$, where $\lambda$ is sampled from a given distribution $\beta$. The output is computed as $\tilde{s}=(K\odot s^{A})\oplus(Q\odot s^{B})$ and the new label $\tilde{y}=y^{A}\oplus y^{B}$. Here, $\odot$ and $\oplus$ represent masking operation and concatenation, respectively. A common choice of $\beta$ distribution is $\beta(2,2)$, which concentrates most values around 0.5. However, radar point clouds can be extremely sparse in certain regions, meaning that applying a single mix ratio uniformly across all points may be suboptimal and limit diversity.
@@
    PolarMix [xiao2022polarmix] randomly divides the field of view into two azimuth regions and swaps all points within the selected angular interval. As shown in Fig. 3(a), a random 90 region is chosen for this exchange, with all points and corresponding labels being swapped. Additionally, points within bounding boxes in $s^{B}$ are copied and rotated around the origin by a randomly selected angle within a predefined range into $\tilde{s}$ to enhance object diversity. Unlike LiDAR sensors, which offer uniform 360 detection, radar point clouds encounter irregularities in their angular distribution due to the radiation pattern, as shown in Fig. 2(a). This leads to significant variations in detection capability across the main lobe and side lobes, especially along the azimuth axis. Accordingly, substantial rotation operations such as PolarMix will disrupt the inherent angular distribution of the point cloud.
@@
    LaserMix [kong2023lasermix] exploits the inclination angle for partitioning by leveraging the distinctive pattern distribution of laser beams. As explained in Fig. 3(b), the points clouds are divided into $N$ equally spaced segments based on the elevation angular interval. An alternating selection of point segments from $s^{A},s^{B}$ is then utilized to construct $\tilde{s}$. However, multi-radar setups are frequently adopted to enhance radar perception, for instance in the Bosch Street dataset [armanious2024bosch], which employs five front-facing radars. As Fig. 2(b) shows, these multiple sensors emit beams from different coordinate systems, so the spatial priors established by laser partitions mentioned in LaserMix [kong2023lasermix] no longer hold. Additionally, radar sensors typically have a lower resolution in the elevation angle and a narrower vertical field of view compared to LiDAR [han2023radar] (Fig. 2(c)). Thus, mixing based on elevation angle or polar coordinates may not be applicable to radar sensors due to their limited elevation performance and the effects of multi-sensor aggregation.
@@
    PillarMix [wu2024pillarmix] directly switches the points and labels within selected pillars to generate mixed samples Compared to LiDAR point clouds, radar point clouds are significantly sparser. Replacing portions of these sparse objects with irrelevant or empty pillars may remove critical information needed for accurate predictions. Consequently, training on such augmented data can degrade performance, making PillarMix less effective than MixUp for these objects. Hence, directly switching entire pillars may be suboptimal for radar point clouds, especially when dealing with sparse objects.
As discussed above, the distinct point distribution patterns in radar point clouds, combined with their inherent sparsity and the use of multi-radar sensor fusion, differ significantly from 360 LiDAR data. Therefore, aforementioned methods developed for LiDAR data often fall short when applied to radar. Later results will reveal that some of these approaches not only underperform compared to general mixed-sample techniques like MixUp [zhang2017mixup] but can even negatively impacting perception performance.
@@ III Approach
[…]
@@ III-A Overview
Inspired by PillarMix [wu2024pillarmix], we propose CAPMix, a class-aware variant that accounts for the sparsity of radar point clouds. The overall methods of both PillarMix and CAPMix are illustrated in Fig. 4. Like PillarMix, CAPMix begins by pillarizing two point clouds $s^{A},s^{B}$ using a specified pillar size $\phi$ (covering an area of $\phi\times\phi$ in Bird’s-Eye View). However, instead of using non-adjacent cross-selection as in PillarMix, all the pillars $P$ are randomly divided into $P_{c},P_{u}]$ based on a ratio $\alpha$. Next, MixUp is applied to each pillar, with the mix ratio independently sampled from a predefined $\beta$ distribution. For $P_{c}$, the mix ratio is class-aware, with pillars assigned a mix ratio from ${\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}},\,{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}},\,\textrm{or}\,\,{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$ based on their object density (sparse, moderate, or dense). To enhance diversity while preserving original information, $\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}$} is also applied to all pillars in $P_{u}$ without considering class labels. As a data augmentation technique, CAPMix is employed exclusively during training, while the inference stage remains identical to the standard process.
@@ III-C Class-aware $\beta$ distributions in $P_{c}$
[…]
@@
6:  $P_{c},\,P_{u}]=\textrm{RandomSelect}([P_{A},\,P_{B}],\,\alpha),$
@@
To ensure the model retains sufficient information for accurate predictions, the class-aware mix ratio is applied only to $P_{c}\,$, rather than all pillars. Applying it universally would result in excessive information loss, particularly for large objects, making them significantly harder for the model to detect. For instance, if $\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}$} were applied to dense objects in $P_{u}$, every part of the object would receive a $\lambda$ value close to 0, leading to the removal of most points. In contrast, using $\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}$} for $P_{u}$ ensures that enough points are preserved, maintaining the integrity of the object and facilitating reliable predictions. Compared to PillarMix [wu2024pillarmix], which retains all points in $P_{u}\,$, applying a moderate mix ratio provides a balanced approach, preserving sufficient information for effective detection while still benefiting from augmentation. The operation can be mathematically expressed as follows:
@@
$\tilde{s}$ $=\{\tilde{p}_{c}\}\oplus\{\tilde{p}_{u}\},$
@@
Our method leverages bounding box labels while operating on radar data pillars, recognizing that radar points can extend beyond the bounding box, particularly for moving objects  [palmer2023ego]. By applying pillar-wise operations, we ensure that these points, along with non-object points, are included in the mixing process. Moreover, we use object labels rather than sparsity to distinguish pillars, ensuring sparse objects are not overshadowed or erased when sharing a pillar with dense objects. This careful handling of sparse points with high priority is crucial for maintaining data integrity.